# Summaries as Centroids for Interpretable and Scalable Text Clustering

### v3

@@ 1 Introduction
Figure 1: Illustration of k-NLPmeans/k-LLMmeans with a single summarization step. First panel shows the text embeddings with stars marking the initial centroids; second shows the partition reached after k-means iterations (a local minimum); third performs the summarization step, each previous cluster is summarized into a textual prototype and re-embedded; final panel runs one more k-means iteration using these summaries as centroids, yielding a qualitatively improved partition.
@@
We consider two families of summarizers: (1) Classical NLP—centroid-based summarization [48], graph-based methods such as TextRank [37], and LSA-style techniques [14]—which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content [71, 50, 28] (yielding k-LLMmeans). In both variants, summaries are recomputed only at scheduled iterations, so the algorithms share the same summary-as-centroid mechanism and differ solely in the summarizer. While methods from (1) and (2) are standard in summarization, our periodic use of summary-derived centroids within the k-means loop is, to our knowledge, novel and central to the method’s interpretability.
@@
In summary, we (i) propose k-NLPmeans and k-LLMmeans, a summary-as-centroid variant of k-means that periodically replaces numeric centroids with textual prototypes re-embedded by the same encoder—an LLM-optional design that, by construction, preserves the standard k-means objective between summary steps; interpretability follows from human-readable centroids and transparent intermediate outputs; (ii) extend the approach to mini-batch k-means for efficient, online clustering of streams; (iii) present a comprehensive empirical study across datasets, embeddings, and summarizers, showing consistent gains over classical baselines and competitiveness with recent LLM-based clustering under a dataset-size–independent LLM budget; and (iv) provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
@@ 2 Preliminaries: k-means for text clustering
Lloyd’s algorithm [35]—the standard heuristic for k-means—alternates between assigning each document embedding $\mathbf{x}_{i}$ to its nearest centroid and recomputing each centroid as the mean of its assigned points. Repeating these two steps for $T$ iterations monotonically reduces the within‑cluster sum of squared distances, steering the procedure toward a (locally) optimal set of centroids. However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima [36, 35]. Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results [5].
@@
We enhance k-means for text clustering by periodically replacing the numerical centroid update with summarization steps that yield textual-summary–based centroids. The procedure (Algorithm 1 in Appendix C) is identical to k-means algorithm except that every $l$ iterations the mean update in (2) is replaced by the embedding of a cluster textual summary. At all other iterations, the standard update is used (see Figure 1 for an illustration with one summarization step). The summarizer can be instantiated with classical, deterministic methods (k-NLPmeans) or with LLM-based summaries (k-LLMmeans), and is not restricted to the specific techniques we discuss—any textual summarization operator can be dropped in.
@@ 3.3 Advantages of our approaches
Over k-means. Periodic summarization steps act as a semantic prototype update that can redirect the search trajectory and reduce sensitivity to initialization. Instead of relying solely on Euclidean means, a textual summary captures contextual cues present in the underlying documents; re-embedding this summary yields centroids that better reflect cluster semantics. In practice, summary-based centroids produce more interpretable and often more semantically coherent—partitions, even when k-means++ seeding is suboptimal. Apart from the summary steps, the procedure follows standard k-means: assignments and numeric updates are unchanged and the usual objective is preserved between summary iterations. Section 5.1 shows that our methods often outperform vanilla k-means across diverse settings.
@@ 4 Mini-batch k-NLPmeans and k-LLMmeans
Mini-batch k-means [56] is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset. This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams—such as those from social media, news, or customer feedback—where data must be clustered incrementally without full dataset access. Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
@@ 5 Static experiments
As traditional clustering baselines, we consider k-means, k-medoids [31], spectral clustering [41], agglomerative clustering [30], and Gaussian Mixture Models (GMM) [15]. We use ground-truth number of clusters and ten different seeds. For advanced LLM-based methods, we compare with results obtained by Feng et al. [20] on ClusterLLM [73], IDAS [13], and LLMEdgeRefine [20]. All k-means based methods are initialized with k-means++. We now provide a summary of the results.
@@ 5.1 Results with static data
Table 1: Average ACC and NMI for four k-LLMmeans variants using GPT-4o, compared against traditional baselines, with text-embedding-3-small embeddings, evaluated on benchmark datasets.
@@
Table 2: Average ACC, NMI, and distdist for k-means and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
@@
Table 3: Number of LLM calls (prompts), average ACC, and average NMI for k-LLMmeans FS-multiple using various LLMs and e5-large embeddings, compared against other LLM-based text clustering methods on three benchmark datasets.
@@
Comparing our approaches with different LLMs and state‑of‑the‑art LLM‑based clustering methods. Table 3 reports the number of LLM calls (prompts) and mean ACC/NMI for k-NLPmeans (LSA-multiple) and k-LLMmeans (FS-multiple) on three benchmarks with e5-large embeddings. For k-LLMmeans we vary the backbone generator—including GPT-3.5—and observe stable performance across LLMs, indicating robustness to the specific model. By design, k-NLPmeans uses zero LLM calls. For context, we also include results from Feng et al. [20]: ClusterLLM (fine-tuned) and IDAS use GPT-3.5 with the same e5-large embeddings, while LLMEdgeRefine uses the stronger Instructor embeddings [29]. Our GPT-3.5 configuration of k-LLMmeans achieves comparable, sometimes slightly lower ACC/NMI, but with much fewer LLM calls (independent of dataset size) and no fine-tuning. k-NLPmeans trails k-LLMmeans slightly but remains competitive without any LLM usage. Overall, our framework offers a favorable quality–cost trade-off: k-NLPmeans for zero-inference-cost deployments, and k-LLMmeans for higher accuracy under a small, fixed summarization budget.
Results with standard deviations are reported in Tables 5, 6, and 7 of Appendix B.1. We also assess sensitivity to the instruction prompt II in k-LLMmeans and to parameter $q$ in k-NLPmeans. As shown in Table 9 (Appendix B.2), performance is generally stable across prompt choices and $q$ values.
@@ 6.1 Results with streaming data
Table 4 reports average ACC and NMI on the yearly StackExchange corpus, comparing mini‑batch k-NLPmeans (LSA-multiple) and mini‑batch k-LLMmeans variants against standard k-means, mini‑batch k-means, and sequential mini‑batch k-means. Mini-batch k-LLMmeans variants achieve the highest ACC and NMI—consistently outperforming baselines, and even surpassing full-dataset k-means. Despite operating in a streaming regime, our approach clusters the entire corpus of 205,943 posts using no more than 3,850 LLM calls, demonstrating a favorable accuracy-per-LLM-token trade-off. The few-shot summarization design keeps prompts short, sidestepping context-window limits while preserving cluster quality. Mini-batch k-NLPmeans also improves over all mini-batch baselines, though it trails the k-LLMmeans variants—likely because its extractive summaries are more sensitive to the noisiness and heterogeneity of community posts than LLM-generated centroids. Overall, these results highlight that our interpretable, mini-batch formulation scales to long sequential streams with strong accuracy and limited LLM usage.
@@ 7 Case study
Interpretation. Our interpretable clustering enables temporal analysis of topic change, since each centroid is a human-readable summary rather than a latent vector. Figure 2 shows this for three major clusters: Image Model Optimization, AI Evolution and Challenges, and Advanced NLP Techniques. Across 2021, these clusters drift from fundamentals toward integration-at-scale, mirroring how tools, models, and deployment pressures matured. Image Model Optimization begins with robustness anxieties (adversaries, occlusion, scale) around mainstream detectors/segmenters, pivots mid-year to “how do I implement this well?” (transfer learning, augmentation, preprocessing), and ends with production hardening (class imbalance, specialized losses, edge efficiency) as strong pretrained models and libraries made setup routine and pushed bottlenecks to data and deployment [8, 32]. AI Evolution and Challenges moves from historical “what is AI becoming?” to curiosity about large generative/multimodal systems—probably sparked by public releases like CLIP/DALL·E and their failure modes—and finishes with system-level synthesis (neuro-symbolic integration, RL control, compute efficiency, safety) as enterprise adoption and policy attention grow [49, 51, 40, 42, 19]. Advanced NLP Techniques tracks the transformer boom: early questions focus on which models and preprocessing to use; mid-year attention shifts to measuring semantics beyond BLEU (e.g., BERTScore, BLEURT) as long-context and domain-shift issues surface; late-year concerns are product-driven—sequence limits, domain vocab, multilingual/cross-modal alignment—reflecting rapid uptake of pretrained encoders/decoders in industry [10, 72, 57, 68]. Practically, these trajectories can power better post categorization, searchability, answer routing, and trend detection. More broadly, they showcase how our mini-batch k-LLMmeans exposes readable centroids enabling end-to-end interpretability for sequential text streams, letting practitioners track topic evolution, attribute causes, and make transparent, timely decisions in dynamic corpora.
@@ 8 Related work
Traditional clustering. Hierarchical methods [30, 6] build tree-structured representations of nested document relationships. Density-based approaches like DBSCAN [18] and graph-based methods detect clusters of arbitrary shapes, while spectral clustering [41] leverages eigen-decomposition to uncover complex structures. Model-based techniques—including Gaussian mixture models [15] and recent neural network frameworks [74, 25, 69, 70, 67]—provide probabilistic clustering formulations. Additionally, topic modeling methods, from probabilistic latent semantic analysis [24] to latent Dirichlet allocation [7], capture word co-occurrence patterns and latent topics. BERTopic [23] offers an embedding–reduction–density pipeline for interpretable topics, while DeepCluster [11] uses clustering to iteratively train encoders in a self-supervised manner (orthogonal to our fixed-embedding setting). LLM-based clustering. Viswanathan et al. [64] employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering. Zhang et al. [73] propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity. Complementary approaches [63, 47] show that context-derived representations capture subtle semantic nuances beyond traditional embeddings. Wang et al. [66] introduce a goal-driven, explainable method that employs natural language descriptions to clarify cluster boundaries, while [13] present IDAS for intent discovery using abstractive summarization. Feng et al. [20] propose LLMEdgeRefine an iterative mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
@@ 9 Discussion
Limitations. Our k-LLMmeans relies on LLM-generated summaries; thus, biases or errors in the underlying model can propagate to cluster prototypes and downstream analyses. Although simple instruction prompts are effective (Appendix B.2), prompt design can still influence outcomes. The few-shot variant assumes that a small set of representative samples captures cluster structure—practical in many settings but potentially limiting for heterogeneous clusters. Regarding potential pretraining exposure, note that the LLM in our pipeline is used solely to summarize texts already assigned to each cluster; it does not inject external labels or unseen content, so any exposure does not confer an unfair advantage. As indicated by suboptimal results on the StackExchange dataset (Table 4), k-NLPmeans may benefit from applying text-cleaning preprocessing prior to use.
@@
We release code to reproduce all results in this paper at https://github.com/jairoadiazr/summaryCentroids (see the README for instructions).

### v4

@@ 1 Introduction
Our goal is not to strictly dominate every clustering algorithm, but to provide a simple novel approach that (i) improves clustering quality over standard k-means–style methods, (ii) yields human-readable centroids, and (iii) scales to large and streaming datasets with limited LLM usage.
Figure 1: Illustration of k-NLPmeans/k-LLMmeans with a single summarization step. First panel shows the text embeddings with stars marking the initial centroids; second shows the partition reached after k-means iterations (a local minimum); third performs the summarization step, each previous cluster is summarized into a textual prototype and re-embedded; final panel runs one more k-means iteration using these summaries as centroids, yielding a qualitatively improved partition. This example is illustrative; in practice, clusters can be multi-topic, and summaries may describe multiple prominent subtopics for a single cluster.
@@
We consider two families of summarizers: (1) Classical NLP such as centroid-based summarization [48], graph-based methods such as TextRank [37], and LSA-style techniques [14], which are fast and deterministic (yielding k-NLPmeans); and (2) LLM-based summaries that can capture richer context and filter secondary or noisy content [71, 50, 28] (yielding k-LLMmeans). In both variants, summaries are recomputed only at scheduled iterations, so the algorithms share the same summary-as-centroid mechanism and differ solely in the summarizer. While methods from (1) and (2) are standard in summarization, our periodic use of summary-derived centroids within the k-means loop is, to our knowledge, novel and central to the method’s interpretability.
@@
In summary, we (i) propose k-NLPmeans and k-LLMmeans, a summary-as-centroid variant of k-means that periodically replaces numeric centroids with textual prototypes re-embedded by the same encoder, an LLM-optional design that, by construction, preserves the standard k-means objective between summary steps; interpretability follows from human-readable centroids and transparent intermediate outputs; (ii) extend the approach to mini-batch k-means for efficient, online clustering of streams; (iii) present a comprehensive empirical study across datasets, embeddings, and summarizers, showing consistent gains over classical baselines and competitiveness with recent LLM-based clustering under a dataset-size–independent LLM budget; and (iv) provide a case study on sequential text streams and release a StackExchange-derived benchmark for evaluating streaming text clustering.
@@ 2 Preliminaries: k-means for text clustering
Lloyd’s algorithm [35], the standard heuristic for k-means, alternates between assigning each document embedding $\mathbf{x}_{i}$ to its nearest centroid and recomputing each centroid as the mean of its assigned points. Repeating these two steps for $T$ iterations monotonically reduces the within‑cluster sum of squared distances, steering the procedure toward a (locally) optimal set of centroids. However, due to its sensitivity to initialization and the non-convex nature of its objective function, k-means does not guarantee convergence to the global optimum and can instead become trapped in local optima [36, 35]. Various strategies, such as k-means++ initialization and multiple restarts, have been proposed to mitigate these issues and improve the likelihood of achieving better clustering results [5].
@@
We enhance k-means for text clustering by periodically replacing the numerical centroid update with summarization steps that yield textual-summary–based centroids. The procedure (Algorithm 1 in Appendix D) is identical to k-means algorithm except that every $l$ iterations the mean update in (2) is replaced by the embedding of a cluster textual summary. At all other iterations, the standard update is used (see Figure 1 for an illustration with one summarization step). The summarizer can be instantiated with classical, deterministic methods (k-NLPmeans) or with LLM-based summaries (k-LLMmeans), and is not restricted to the specific techniques we discuss, any textual summarization operator can be dropped in.
@@ 3.3 Advantages of our approaches
Over k-means. Periodic summarization steps act as a semantic prototype update that can redirect the search trajectory and reduce sensitivity to initialization. Instead of relying solely on Euclidean means, a textual summary captures contextual cues present in the underlying documents; re-embedding this summary yields centroids that better reflect cluster semantics. In practice, summary-based centroids produce more interpretable and often more semantically coherent partitions, even when k-means++ seeding is suboptimal. Apart from the summary steps, the procedure follows standard k-means: assignments and numeric updates are unchanged and the usual objective is preserved between summary iterations. Section 5.1 shows that our methods often outperform vanilla k-means across diverse settings.
@@ 4 Mini-batch k-NLPmeans and k-LLMmeans
Mini-batch k-means [56] is an efficient strategy for large-scale text clustering that processes small, randomly sampled mini-batches instead of the full dataset. This approach substantially reduces memory usage and computational cost, making it well suited for continuously generated text streams, such as those from social media, news, or customer feedback; where data must be clustered incrementally without full dataset access. Mini-batch k-means exhibits convergence properties comparable to standard k-means while offering superior scalability.
@@ 5 Static experiments
As traditional clustering baselines, we consider k-means, k-medoids [31], spectral clustering [41], agglomerative clustering [30], and Gaussian Mixture Models (GMM) [15]. We use ground-truth number of clusters and ten different seeds. We also include BERTopic [23] as a strong topic-modeling baseline. In addition, we report results obtained by applying our proposed methods to the document embeddings produced by BERTopic. For advanced LLM-based methods, we compare with results obtained by Feng et al. [20] on ClusterLLM [73], IDAS [13], and LLMEdgeRefine [20]. All k-means based methods are initialized with k-means++.
We evaluate performance using clustering accuracy (ACC), which measures the best one-to-one alignment between predicted clusters and gold labels, and Normalized Mutual Information (NMI), which quantifies the mutual information between them normalized to $\left[0,1\right]$ (See Appendix C for details on the metrics). We now provide a summary of the results.
@@ 5.1 Results with static data
Table 1: Average ACC and NMI for k-NLPmeans and k-LLMmeans variants using GPT-4o, compared against traditional baselines, BERTopic, and our k-NLPmeans LSA-multiple and k-LLMmeans FS-multiple variants applied to BERTopic embeddings, using text-embedding-3-small embeddings on benchmark datasets.
@@
Table 2: Average ACC, NMI, and distdist for k-means, k-NLPmeans LSA-multiple and k-LLMmeans FS-multiple, evaluated on three datasets using four different embedding models.
@@
Table 3: Number of LLM calls (prompts), average ACC, and average NMI for k-NLPmeans (LSA-multiple) and k-LLMmeans (FS-multiple) using various LLMs with e5-large embeddings, compared against BERTopic, our variants applied to BERTopic embeddings, and other state-of-the-art LLM-based clustering methods on three benchmark datasets.
@@
We also compare with BERTopic and apply our methods to its document embeddings. While BERTopic is strong on its own, adding our variants consistently improves its performance. Across all datasets, the best results come from either our stand-alone methods or their BERTopic-based versions.
@@
Comparing our approaches with different LLMs and state‑of‑the‑art LLM‑based clustering methods. Table 3 reports the number of LLM calls (prompts) and mean ACC/NMI for k-NLPmeans (LSA-multiple) and k-LLMmeans (FS-multiple) on three benchmarks with e5-large embeddings. For k-LLMmeans we vary the LLM generator, including GPT-3.5, and observe stable performance across LLMs, indicating robustness to the specific model. By design, k-NLPmeans uses zero LLM calls. For context, we also include results from Feng et al. [20]: ClusterLLM (fine-tuned) and IDAS use GPT-3.5 with the same e5-large embeddings, while LLMEdgeRefine uses the stronger Instructor embeddings [29]. Our GPT-3.5 configuration of k-LLMmeans achieves comparable, sometimes slightly lower ACC/NMI, but with much fewer LLM calls (independent of dataset size) and no fine-tuning. k-NLPmeans trails k-LLMmeans slightly but remains competitive without any LLM usage. We also evaluate BERTopic and apply our variants to its embeddings. In this setting, our methods achieve results extremely close to the best LLM-based models while requiring none or only a small number of LLM calls. Overall, our framework offers a favorable quality–cost trade-off: k-NLPmeans for zero-inference-cost deployments, and k-LLMmeans for higher accuracy under a small, fixed summarization budget.
Additional results and experiments.
Results with standard deviations are reported in Tables 5, 6, and 7 of Appendix B.1. We also assess sensitivity to the selection of the number of clusters $k$ in Appendix B.2.1. Table 9 shows robustness when the number of clusters mildly differs from the ground truth clusters, consistently better than vanilla k-means. We also evaluate sensitivity to the instruction prompt $I$ in k-LLMmeans and to parameter $q$ in k-NLPmeans in Appendix B.2.2. As shown in Table 10, performance remains generally stable across prompt choices and $q$ values. Table 11 in Appendix B.2.3 also compares several sampling strategies for the few-shot (FS) version of k-LLMmeans, showing that our k-means++ strategy has consistent performance.
Summary of static experimental results
Across our experiments, our methods (i) improve over classical and topic-modeling baselines (Table 1); (ii) outperforms k-means across embeddings (Table 2); and (iii) approaches strong LLM-based clustering methods at much lower LLM cost (Table 3).
@@ 6.1 Results with streaming data
Table 4 reports average ACC and NMI on the yearly StackExchange corpus, comparing mini‑batch k-NLPmeans (LSA-multiple) and mini‑batch k-LLMmeans variants against standard k-means, mini‑batch k-means, and sequential mini‑batch k-means. Mini-batch k-LLMmeans variants achieve the highest ACC and NMI, consistently outperforming baselines, and even surpassing full-dataset k-means. Despite operating in a streaming regime, our approach clusters the entire corpus of 205,943 posts using no more than 3,850 LLM calls, demonstrating a favorable accuracy-per-LLM-token trade-off. The few-shot summarization design keeps prompts short, sidestepping context-window limits while preserving cluster quality. Mini-batch k-NLPmeans also improves over all mini-batch baselines, though it trails the k-LLMmeans variants, likely because its extractive summaries are more sensitive to the noisiness and heterogeneity of community posts than LLM-generated centroids. Overall, these results highlight that our interpretable, mini-batch formulation scales to long sequential streams with strong accuracy and limited LLM usage.
@@ 7 Case study
Interpretation. Our interpretable clustering enables temporal analysis of topic change, since each centroid is a human-readable summary rather than a latent vector. Figure 2 shows this for three major clusters: Image Model Optimization, AI Evolution and Challenges, and Advanced NLP Techniques. Across 2021, these clusters drift from fundamentals toward integration-at-scale, mirroring how tools, models, and deployment pressures matured. Image Model Optimization begins with robustness anxieties (adversaries, occlusion, scale) around mainstream detectors/segmenters, pivots mid-year to “how do I implement this well?” (transfer learning, augmentation, preprocessing), and ends with production hardening (class imbalance, specialized losses, edge efficiency) as strong pretrained models and libraries made setup routine and pushed bottlenecks to data and deployment [8, 32]. AI Evolution and Challenges moves from historical “what is AI becoming?” to curiosity about large generative/multimodal systems, probably sparked by public releases like CLIP/DALL·E and their failure modes, and finishes with system-level synthesis (neuro-symbolic integration, RL control, compute efficiency, safety) as enterprise adoption and policy attention grow [49, 51, 40, 42, 19]. Advanced NLP Techniques tracks the transformer boom: early questions focus on which models and preprocessing to use; mid-year attention shifts to measuring semantics beyond BLEU (e.g., BERTScore, BLEURT) as long-context and domain-shift issues surface; late-year concerns are product-driven, sequence limits, domain vocab, multilingual/cross-modal alignment; reflecting rapid uptake of pretrained encoders/decoders in industry [10, 72, 57, 68]. Practically, these trajectories can power better post categorization, searchability, answer routing, and trend detection. More broadly, they showcase how our mini-batch k-LLMmeans exposes readable centroids enabling end-to-end interpretability for sequential text streams, letting practitioners track topic evolution, attribute causes, and make transparent, timely decisions in dynamic corpora.
@@ 8 Related work
Traditional clustering. Hierarchical methods [30, 6] build tree-structured representations of nested document relationships. Density-based approaches like DBSCAN [18] and graph-based methods detect clusters of arbitrary shapes, while spectral clustering [41] leverages eigen-decomposition to uncover complex structures. Model-based techniques; including Gaussian mixture models [15] and recent neural network frameworks [74, 25, 69, 70, 67]; provide probabilistic clustering formulations. Additionally, topic modeling methods, from probabilistic latent semantic analysis [24] to latent Dirichlet allocation [7], capture word co-occurrence patterns and latent topics. BERTopic [23] offers an embedding–reduction–density pipeline for interpretable topics, while DeepCluster [11] uses clustering to iteratively train encoders in a self-supervised manner (orthogonal to our fixed-embedding setting).
LLM-based clustering. Viswanathan et al. [64] employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering. Zhang et al. [73] propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity. Complementary approaches [63, 47] show that context-derived representations capture subtle semantic nuances beyond traditional embeddings. Wang et al. [66] introduce a goal-driven, explainable method that employs natural language descriptions to clarify cluster boundaries, while [13] present IDAS for intent discovery using abstractive summarization. Feng et al. [20] propose LLMEdgeRefine an iterative mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
@@ 9 Discussion
Limitations. Our k-LLMmeans relies on LLM-generated summaries; thus, biases or errors in the underlying model can propagate to cluster prototypes and downstream analyses. Although simple instruction prompts are effective (Appendix B.2), prompt design can still influence outcomes. The few-shot variant assumes that a small set of representative samples captures cluster structure, practical in many settings but potentially limiting for heterogeneous clusters. Regarding potential pretraining exposure, note that the LLM in our pipeline is used solely to summarize texts already assigned to each cluster; it does not inject external labels or unseen content, so any exposure does not confer an unfair advantage. As indicated by suboptimal results on the StackExchange dataset (Table 4), k-NLPmeans may benefit from applying text-cleaning preprocessing prior to use. Finally, we illustrate qualitative failure cases in Appendix E.
@@
We release anonymized code to reproduce all results in this paper at https://anonymous.4open.science/r/summaryCentroids-877C/ (see the README for instructions).